OpenStudy (anonymous):

A certain liquid has a vapor pressure of 92.0 Torr at 23.0 °C and 313.0 Torr at 45.0 °C. Calculate the value of ΔH°vap for this liquid. Calculate the normal boiling point of this liquid. How would I do this? Please help

OpenStudy (anonymous):

@Rohitkhanna

OpenStudy (petiteme):

Should you use Clausius-Clapeyron equation?

OpenStudy (anonymous):

do I just solve for T2 and use the same values?

OpenStudy (petiteme):

Well, let the normal boiling point be T2

OpenStudy (anonymous):

wouldn't that just give me the T that is given in the question ?

OpenStudy (petiteme):

Nope. Sorry :/ I suck at explaining chemistry

OpenStudy (anonymous):

i just get a very small number

OpenStudy (petiteme):

What did you get for ΔH°vap?

OpenStudy (anonymous):

im not good at chemistry but let's try

OpenStudy (anonymous):

43.55 kj/mol

OpenStudy (petiteme):

I used \[\log\frac{ P2 }{ P1 }=\frac{ ΔH }{2.303R }*(\frac{ 1? }{ T1 }-\frac{ 1 }{ T2 })\]

OpenStudy (petiteme):

Erase the ?

OpenStudy (anonymous):

what is v?

OpenStudy (petiteme):

I got 49.78kj

OpenStudy (petiteme):

Lol. I don't know if i'm right. Ignore me :))

OpenStudy (anonymous):

volume i dont know if that's right , it's been i while i dont see chemistry

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

sing those number? i dont have V so i dont think i can use that

OpenStudy (petiteme):

You saw my equation?

OpenStudy (anonymous):

yea

OpenStudy (petiteme):

Have you tried it?

OpenStudy (anonymous):

yea...and i got 43.55...its ok..i will do the question tmr...im just really tired now

OpenStudy (petiteme):

Okies. Just close this question ^_^